# The LLM Language Network: A Neuroscientific Approach for Identifying Causally Task-Relevant Units

### v1

@@ 4 Experimental Setup
To validate the model language units’ alignment to the human language network, we employ two approaches: i) investigating whether the model units can replicate landmark neuroscience studies that qualitatively describe the response profiles observed in the human language regions, and ii) quantitatively testing the alignment of language units with brain responses from the human language network. For the first approach, we closely follow the analyses in Fedorenko et al. (2010) and Shain et al. (2024), using the same set of experimental conditions which are commonly used in neuroimaging studies examining lexical and syntactic processing. For the second approach, we measure how well the language units can predict brain activity in the human language network. To do so, we utilize the Tuckute2024 Tuckute et al. (2024b) and Pereira2018 Pereira et al. (2018) benchmarks on the Brain-Score platform Schrimpf et al. (2018, 2020b). Tuckute2024 consists of brain responses from 5 participants who each read 1,000 linguistically diverse sentences, while Pereira2018 consists of 15 subjects reading short passages presented one sentence at a time spanning various different topics. See Appendix E for more details about the datasets.
@@ 7 Similarity Between the Language Network in LLMs and Brains
Beyond qualitative alignment between LLM language units and brains, we investigate the quantitative alignment to brain data. Following standard practice in measuring brain alignment, we train a ridge regression to predict brain activity from model representations, using the same input stimuli presented to human participants in neuroimaging studies Naselaris et al. (2011); Schrimpf et al. (2021). We then measure the Pearson correlation between the predicted brain activations and the actual brain activations of human participants on a held-out set. This process is repeated over 10 cross-validation splits, and we report the average (mean) Pearson correlation as our final result which we here refer to as Brain-Score Schrimpf et al. (2018, 2020b). Figure 5 shows the average raw correlation when using {0.03125, 0.0625, 0.125, 0.25, 0.5, 1}of model units to predict brain activity for two neural datasets Pereira et al. (2018); Tuckute et al. (2024b). This analysis is repeated for the most language selective units, and for three sets of randomly sampled units for each of the 10 models. See Appendix E for more results and dataset details.
@@ 9 Discussion
Our paper adds to the growing body of research that uses neuroscience tools to interpret machine learning models Schrimpf et al. (2020a); Zador et al. (2023); Tuckute et al. (2024a). Specifically, our work takes a step towards analyzing LLM units that are causally useful within a given system, providing a more stringent notion of functional correspondence Cao (2022); Cao and Yamins (2024); Mahon (2023); Prince et al. . The consistency between the causally important language units in LLMs and the human brain may suggest that computations, beyond representations, could be shared between these two systems. This prompts an intriguing question: do specialized subsystems, such as the language network, always emerge as a consequence of optimizing for next-word prediction, and is such a simple objective the driver of specialization in the brain? Exploring this connection further could shed light on how cognitive processes evolve from such predictive mechanisms.
@@ D.1 Language Results
Tables 3 and 4 display results for the 10 models tested on three language benchmarks—SyntaxGym, BLiMP, and GLUE—along with the average performance across these benchmarks. Each model’s performance is shown initially without ablation, followed by ablations of the top-0.125, 0.25, 0.5, 1% language-selective units, and then with randomly sampled units at the same percentages. The performance changes in Figure 3 can be reproduced by subtracting post-ablation results from the baseline (0%) for both language-selective and random unit ablations. Results with random units are averaged across three random seeds.
@@ D.3 Theory of Mind Results
Appendix E Brain-Score Datasets
@@
This dataset consists of 5 participants reading 1000 6-word sentences presented on one sentence at a time for 2s. BOLD responses from voxels in the language network were averaged within each participant and then across participants to yield an overall average language network response to each sentence. The stimuli used span a large part of the linguistic space, enabling model-brain comparisons across a wide range of single sentences. Sentence presentation order was randomized across participants. The averaging of sentences across participants effectively minimizes the effect of temporal autocorrelation/context in this dataset. In combination with the diversity in linguistic materials, this dataset presents a particularly challenging dataset for model evaluation. The noise ceiling for Tuckute2024 is $r=0.56$, see Tuckute et al. (2024b) for more details.
@@
Model Ablation Units Level 1 Level 2 Level 3 Level 4 Level 5 Average

### v2

@@ 4 Experimental Setup
To validate the model language units’ alignment to the human language network, we employ two approaches: i) investigating whether the model units can replicate landmark neuroscience studies that qualitatively describe the response profiles observed in the human language regions, and ii) quantitatively testing the alignment of language units with brain responses from the human language network. For the first approach, we closely follow the analyses in Fedorenko et al. (2010) and Shain et al. (2024), using the same set of experimental conditions which are commonly used in neuroimaging studies examining lexical and syntactic processing. For the second approach, we measure how well the language units can predict brain activity in the human language network. To do so, we utilize the Tuckute2024 Tuckute et al. (2024b) and Pereira2018 Pereira et al. (2018) benchmarks on the Brain-Score platform Schrimpf et al. (2018, 2020). Tuckute2024 consists of brain responses from 5 participants who each read 1,000 linguistically diverse sentences, while Pereira2018 consists of 15 subjects reading short passages presented one sentence at a time spanning various different topics. See Appendix F for more details about the datasets.
@@ 7 Similarity Between the Language Network in LLMs and Brains
Beyond qualitative alignment between LLM language units and brains, we investigate the quantitative alignment to brain data. Following standard practice in measuring brain alignment, we train a ridge regression to predict brain activity from model representations, using the same input stimuli presented to human participants in neuroimaging studies Naselaris et al. (2011); Schrimpf et al. (2021). We then measure the Pearson correlation between the predicted brain activations and the actual brain activations of human participants on a held-out set. This process is repeated over 10 cross-validation splits, and we report the average (mean) Pearson correlation as our final result which we here refer to as Brain-Score Schrimpf et al. (2018, 2020). Figure 5 shows the average raw correlation when using {0.03125, 0.0625, 0.125, 0.25, 0.5, 1}of model units to predict brain activity for two neural datasets Pereira et al. (2018); Tuckute et al. (2024b). This analysis is repeated for the most language selective units, and for three sets of randomly sampled units for each of the 10 models. See Appendix E for more statistical tests and Appendix F for more dataset details.
@@ 9 Discussion
Our paper adds to the growing body of research that uses neuroscience tools to interpret machine learning models Schrimpf et al. (2020, 2021); Zador et al. (2023); Tuckute et al. (2024a). Specifically, our work takes a step towards analyzing LLM units that are causally useful within a given system, providing a more stringent notion of functional correspondence Cao (2022); Cao and Yamins (2024); Mahon (2023); Prince et al. (2024). The consistency between the causally important language units in LLMs and the human brain may suggest that computations, beyond representations, could be shared between these two systems. This prompts an intriguing question: do specialized subsystems, such as the language network, always emerge as a consequence of optimizing for next-word prediction, and is such a simple objective the driver of specialization in the brain? Exploring this connection further could shed light on how cognitive processes evolve from such predictive mechanisms.
@@ D.1 Language Results
Tables 6 and 7 display results for the 10 models tested on three language benchmarks—SyntaxGym, BLiMP, and GLUE—along with the average performance across these benchmarks. Each model’s performance is shown initially without ablation, followed by ablations of the top-0.125, 0.25, 0.5, 1% language-selective units, and then with randomly sampled units at the same percentages. The performance changes in Figure 3 can be reproduced by subtracting post-ablation results from the baseline (0%) for both language-selective and random unit ablations. Results with random units are averaged across three random seeds.
@@ D.3 Theory of Mind Results
Appendix E More Brain Alignment Statistical Tests
In Section 7, we performed Welch’s t-test to demonstrate that units from the language network are significantly more brain-aligned than three sets of randomly sampled units from the model, particularly when sampling a small number of units. Here, we conduct the Shapiro-Wilk test to verify that each distribution follows a normal distribution, as Welch’s t-test assumes normality in the compared distributions. Tables 3 and 4 present the test statistics and p-values for the brain alignment results across models, comparing both language and random units at each percentage and for each dataset. These results confirm that the distributions are indeed normal. A p-value greater than 0.05 indicates normality, while values below this threshold suggest deviation from normality. Notably, the only cases where the p-value falls below 0.05—indicating non-normal distributions—are for the 0.5% and 1% conditions in the Tuckute2024 dataset, where no significant difference was observed.
Percentage Language Units Random Units
0.03125% (0.902, 0.233) (0.971, 0.565)
0.0625% (0.957, 0.755) (0.939, 0.085)
0.125% (0.955, 0.722) (0.981, 0.853)
0.25% (0.933, 0.475) (0.962, 0.345)
0.5% (0.945, 0.609) (0.974, 0.658)
1% (0.945, 0.612) (0.970, 0.551)
Table 3: Shapiro-Wilk test (statistics and p-values) for brain alignment distributions across models. The test is conducted separately for language units and randomly sampled units at each percentage level for the Pereira2018 dataset.
Percentage Language Units Random Units
0.03125% (0.948, 0.641) (0.9511, 0.181)
0.0625% (0.962, 0.802) (0.9507, 0.176)
0.125% (0.973, 0.915) (0.9810, 0.852)
0.25% (0.914, 0.309) (0.9592, 0.296)
0.5% (0.829, 0.032) (0.9693, 0.519)
1% (0.825, 0.029) (0.9725, 0.608)
Table 4: Shapiro-Wilk test (statistics and p-values) for brain alignment distributions across models. The test is conducted separately for language units and randomly sampled units at each percentage level for the Tuckute2024 dataset.
We also conducted a permutation test, a non-parametric statistical method that does not require the assumption of normality. This method involves randomly shuffling data labels across 10,000 permutations to generate a null distribution of the test statistic. By comparing the observed test statistic to this null distribution, we evaluated the statistical significance of our results. The findings from the permutation test confirmed the significance of our results, as shown in Table 5.
Percentage Pereira2018 Tuckute2024
0.03125% 0.001 0.004
0.0625% 0.004 0.013
0.125% 0.138 0.000
0.25% 0.548 0.013
0.5% 0.195 0.169
1% 0.084 0.458
Table 5: Permutation test p-values assessing the statistical significance of brain alignment differences on both datasets. The test was conducted by randomly shuffling data labels across 10,000 permutations to generate a null distribution of the test statistic. The observed test statistic was then compared to this null distribution to compute the p-values. Lower p-values indicate stronger evidence against the null hypothesis, confirming the robustness of our findings.
Appendix F Brain-Score Datasets
@@
This dataset consists of 5 participants reading 1000 6-word sentences presented one sentence at a time for 2s. BOLD responses from voxels in the language network were averaged within each participant and then across participants to yield an overall average language network response to each sentence. The stimuli used span a large part of the linguistic space, enabling model-brain comparisons across a wide range of single sentences. Sentence presentation order was randomized across participants. The averaging of sentences across participants effectively minimizes the effect of temporal autocorrelation/context in this dataset. In combination with the diversity in linguistic materials, this dataset presents a particularly challenging dataset for model evaluation. The noise ceiling for Tuckute2024 is $r=0.56$, see Tuckute et al. (2024b) for more details.
@@
Model Ablation Level 1 Level 2 Level 3 Level 4 Level 5 Average